Cobalt reacts with excess aqueous hydrochloric acid to produce hydrogen. (a) Write a balanced chemical equation for the reaction, using the smallest integer coefficients possible. Write states in brackets. (Hint: CoCl2 is the second product.) Co + 2HCl --> 2H + CoCl2 (Is this formula correct?) (b) Calculate the mass of pure Co that will furnish 22.9 L of hydrogen at a pressure of 0.785 atm and a temperature of 27.6°C. = g
Answer #1

a) balanced chemical reaction between cobalt and HCl is given bellow

ans Co(s) + 2HCl(g) H2(g) + CoCl2(s)

b) volume of H2 gas produced = 22.9 liter , pressure = 0.785 atm , temperature = 27.60 C

use ideal gas equation to find no of moles of H2 gas produced

Ideal gas equation

PV = nRT             where, P = atm pressure= 0.785 atm,

V = volume in Liter = 22.9 L

n = number of mole = ?

R = 0.08205L atm mol-1 K-1 =Proportionality constant = gas constant,

T = Temperature in K = 27.60C = 273.15+ 27.6= 300.75 K

We can write ideal gas equation

n = PV/RT

Substitute the value

n = (0.785 22.9)/(0.08205 300.75) = 0.728 moles

0.728 moles of H2 gas are produced

according to balanced chemical equation given in part a for production of 1 mole of hydrogen gas requires 1 mole cobalt therefore for production of of 0.728 moles of hydrogen gas will require 0.728 mole of cobalt

no of moles of Co required = 0.728 moles , molar mass of Co = 58.93 gram/mole

mass in gram = no of moles x molar mass , substitute values to find grams of cobalt required

mass of cobalt required = 0.728 x 58.93 = 42.9 gram

ans 42.9 gram of pure cobalt will furnish 22.9 liter of hydrogen gas at 0.785 atm pressure and 27.60 C temperature